A 1.25×10^-4 M solution of the anti-inflammatory drug naproxen has a pH = 4.2. Find the Ka and pKa of naproxen.
Answer #1


here the pH = 4.2

pH = -log[H+]

=> [H+] = 6.31*10-5 M

Naproxen is an carboxylic acid and it is taken as HA

     HA -------------------> H+ + A-

I   0.000125                    0     0

C      -x                           +x   +x

E (0.000125-x)                x    x

Ka = [H+][A-] / [HA]

Ka = (6.31*10-5 ) (6.31*10-5 ) / (0.000125-6.31*10-5)

Ka = 6.43*10-5

pKa = -log[Ka]

pKa = -log[6.43*10-5 ]

pKa = 4.1916